OpenStudy (anonymous):

Which function has real zeros at x = 3 and x = 7? f(x) = x2 + 4x – 21 f(x) = x2 – 4x – 21 f(x) = x2 – 10x + 21 f(x) = x2 – 10x – 21

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@iGreen @iambatman

OpenStudy (anonymous):

x = 3 x-3 = 0 x=7 x-7 =0 (x-3)(x-7)=0 find function

OpenStudy (sleepyjess):

Break it down, if the zeros are 3 and 7, the binomials are going to be (x-3)*(x-7). Do you know how to multiply these?